### v1

# ReGATE: Learning Faster and Better with Fewer Tokens in MLLMs

Abstract
The computational cost of training multimodal large language models (MLLMs) rapidly increases with the number of tokens involved. Existing efficiency methods primarily target inference and rely on token reduction or merging, offering limited benefit during training. In this paper, we propose ReGATE (Reference-Guided Adaptive Token Elision), an adaptive token pruning method for accelerating MLLM training. Specifically, ReGATE adopts a teacher-student framework in which the MLLM being trained serves as the student, and a frozen reference large language model (LLM) acts as the teacher. The teacher computes per-token reference losses, which are combined with an exponential moving average (EMA) of the student’s own difficulty scores. This adaptive difficulty-based scoring enables the selective processing of crucial tokens while bypassing less informative ones in the forward pass, significantly reducing computational overhead. Experiments demonstrate that ReGATE, when applied to VideoLLaMA2, matches the peak accuracy of standard training on MVBench up to 2 faster, using only 35% of the tokens. With additional training, it even surpasses the baseline on several multimodal benchmarks, all while reducing the total token count by over 41%. Code and models will be released soon.
ReGATE: Learning Faster and Better with Fewer Tokens in MLLMs
Chaoyu Li, Yogesh Kulkarni, Pooyan Fazli Arizona State University, Arizona, USA {chaoyuli, ykulka10, pooyan}@asu.edu
1 Introduction
Multimodal large language models (MLLMs) face significant challenges due to the high computational cost of training. A key bottleneck is the self-attention mechanism, whose complexity grows quadratically with input sequence length Vaswani et al. (2017). This problem is amplified in video tasks, where frames are tokenized into extremely long sequences. Consequently, training MLLMs on large-scale instructional datasets demands substantial computing resources, limiting accessibility and slowing progress in the field.
Several strategies have been proposed to speed up inference in MLLMs, including static token pruning Arif et al. (2025) and token merging Chen et al. (2024). However, reducing the high cost of training remains a more complex and less explored challenge. In the unimodal text domain, recent work such as RHO-1 has introduced learnable token pruning techniques that improve training efficiency Lin et al. (2024c). Yet, these training-time acceleration methods have not been extended to large multimodal models. Earlier attempts to improve visual processing efficiency, typically targeting standard vision transformers (Akbari et al., 2021) or early video-language models (Lei et al., 2021), have relied on heuristic approaches such as random token dropping. These methods fall short in modern MLLMs, as they fail to capture the subtle and often unintuitive cross-modal importance of tokens, particularly in video, where information is both dense and temporally distributed. As a result, such methods risk discarding important visual or semantic content, which can lead to unstable training and weaker multimodal understanding.
To address this challenge, we introduce ReGATE (Reference-Guided Adaptive Token Elision), a framework designed to accelerate the training of MLLMs. ReGATE adopts a teacher-student architecture, where the student is the multimodal model being trained, and the teacher is a frozen, text-only version of the same LLM backbone. This setup enables ReGATE to dynamically identify and retain the most informative tokens during training by combining two complementary signals. First, it assesses whether a token requires visual grounding by checking if the text-only teacher can accurately predict it from the prompt alone. Second, it evaluates the student model’s learning progress using an exponential moving average (EMA) of token-wise historical losses. By integrating these signals, ReGATE allocates computation to the subset of tokens that are both critical for multimodal understanding and remain challenging for the model to learn.
To summarize, our contributions are threefold:
We introduce ReGATE, an adaptive token pruning method for accelerating MLLM training. ReGATE leverages a text-only reference teacher model and the student’s historical token difficulty to dynamically identify and retain visually essential tokens, without introducing any additional trainable parameters.
We show that the model-agnostic ReGATE integrates seamlessly into existing MLLMs, requiring no architectural changes, making it easy to adopt.
Extensive experiments on image and video benchmarks demonstrate ReGATE’s broad applicability and efficiency. Notably, on the challenging MVBench benchmark, ReGATE, when applied to VideoLLaMA2, matches the baseline’s peak accuracy in just 16.0 hours (compared to 32.4 hours for standard fine-tuning) while processing only 29.3 million tokens, a 65% reduction from the baseline’s 83.8 million (Figure 1).
2 Related Work
2.1 Token Compression for Fast Inference
Most existing work in the literature focus on accelerating inference, not training. Inference-time sparsity methods have shown that many tokens can be removed or merged with minimal impact on accuracy. In vision transformers, Dynamic Token PruningTang et al. (2023) halts processing of “easy” tokens layer by layer, reducing FLOPs by 20–35% on semantic segmentation tasks without degrading performance. For video LLMs, DyCokeTao et al. (2025) dynamically compresses spatial-temporal tokens during inference, achieving up to 2 speed-ups while keeping model weights frozen. Moving from pruning to aggregation, Importance-Based Token Merging Wu et al. (2025) merges highly similar tokens rather than dropping them, maintaining performance on long-video benchmarks while delivering 1.5 faster inference. However, all these methods operate after training is complete. During training, the full token sequence is still processed in every forward and backward pass, leaving the computational cost of training mainly unaddressed.
2.2 Token Compression for Fast Training
Only a few studies have explored token compression during training, rather than just at inference. In text-only language models, RHO-1 Lin et al. (2024c) ranks tokens with a reference model and backpropagates only through the most difficult subset, reducing pre-training tokens by 50% while improving accuracy. For MLLMs, LaVi Yue et al. (2025) avoids processing long visual sequences by injecting vision-conditioned deltas—small, token-specific offsets derived from the visual input—into layer norms, eliminating most visual tokens, but this requires a specialized modulation pathway that must be trained from scratch. LLaVA-Meteor li et al. (2025) introduces a flash-fusion module and a dual-expert scorer that prunes 75–95% visual tokens during instruction tuning but adds extra parameters and targets only vision tokens. In contrast, ReGATE uniquely combines two complementary difficulty signals: a static, cross-modal reference loss from a frozen text-only teacher that identifies tokens requiring visual grounding, and a dynamic learning signal based on the student model’s own token-wise loss tracked via an exponential moving average (EMA). This fusion of global and local difficulty enables a highly adaptive, parameter-free sparsity mechanism that gates both text and vision tokens, without modifying the underlying model architecture.
2.3 Teacher-Student Distillation for MLLMs
Most distillation approaches for MLLMs mainly focus on parameter compression. A systematic study Xu et al. (2024) shows that jointly aligning tokens and logits helps a smaller student model inherit visual grounding from a larger teacher model. Similarly, methods like DIME-FM Sun et al. (2023) show how cross-modal features can be transferred even from unpaired data. A more recent approach, MaskedKD Son et al. (2024), improves efficiency by masking a portion of the image patch tokens fed to the teacher based on the student’s attention scores. This strategy saves up to 50% of the teacher’s FLOPs without reducing student accuracy. However, MaskedKD only sparsifies the teacher’s computation and still requires backpropagation through all student tokens. In contrast, ReGATE introduces a fundamentally different approach by redefining the teacher’s role in distillation. It uses the teacher’s per-token loss to decide which tokens the student should process during each forward and backward pass. Instead of focusing on compressing the model itself, ReGATE targets compressing the computation path. This novel paradigm provides on-the-fly, modality-agnostic sparsity that optimizes the training process without changing the student’s backbone architecture.
3 ReGATE
We introduce ReGATE, a method that speeds up the training of MLLMs by selectively allocating computational resources only to tokens that truly require visual information. The key insight is that not all tokens in a multimodal sequence depend equally on visual context: some can be accurately predicted from text alone, while others need cross-modal grounding. To capture this, ReGATE uses a teacher-student framework. The student is the main MLLM being trained. The teacher is a reference model created by taking the student’s LLM backbone, removing its visual components (the visual encoder and projector), and freezing its weights. This results in a pure text-only LLM that acts as a fixed expert to estimate the degree to which each token depends on visual input. Given a batch of input sequences containing both text and visual tokens, we generate a binary mask that determines which token positions should be actively computed and which can be skipped. This section explains how we calculate per-token difficulty scores using the frozen text-only teacher combined with the student’s own training history, how we dynamically adjust the fraction of tokens retained during training, and how we apply the resulting mask within the transformer decoder.
3.1 Difficulty Score Formulation
Let denote the token sequence in sample , including both text tokens and special visual tokens (e.g., <image> or <video> tokens representing visual content). To compute the reference loss, we construct a modified sequence by replacing the actual visual tokens with placeholder tokens (typically the padding token <pad>), ensuring the sequence length remains identical to the original multimodal input fed to the MLLM’s backbone LLM. Our reference model is a pure text-only LLM obtained by removing the visual encoder and projector from the MLLM backbone, thus incapable of processing any visual content. By feeding the constructed placeholder sequence to the reference model in evaluation mode, we compute the per-token negative log-likelihood:
A low value of indicates that the teacher can predict based on the textual context alone, whereas a high value signals that multimodal information is needed to predict the token. In parallel, we monitor how difficult each token has been for the student across training updates. For every training sample and token position , we maintain a running difficulty buffer updated as an exponential moving average (EMA) of the student’s cross-entropy loss:
where is the current cross-entropy loss of the student model at token position , and controls the smoothing of the EMA. A higher value of indicates that token in sample has consistently posed difficulties during training. We then combine the reference loss and the student’s historical difficulty into a unified difficulty score for each token:
where balances these two signals. Tokens with a higher combined difficulty, , either consistently challenge the student model or genuinely require visual context, and thus are prioritized during the training updates. Note that this combined difficulty evaluation is performed exclusively on output tokens (labels), as these tokens directly influence the training process through backpropagation.
3.2 Dual-cycle Sparsity Schedule
We employ a deterministic schedule to determine the fraction of tokens kept at each training step. Our schedule repeats every steps. In the first steps of each cycle, we keep all tokens (i.e., ) to allow the model to stabilize. In the remaining steps, we retain only a fixed proportion of the tokens. Formally, if denotes the global training step, we have:
3.3 Dynamic Token Gating
For each sample , we identify the indices of valid tokens excluding padding and special markers. Let denote those indices and . We compute the combined difficulty for using Equation (3) and select the top tokens. The resulting binary mask is set to one for retained tokens and zero otherwise. We always retain all special visual tokens (e.g., those corresponding to a frame or image) regardless of their difficulty to preserve multimodal information.
Because the difficulty buffer is updated after every epoch, the set of selected positions adapts throughout training: tokens that become easy for the student are gradually deprioritised, while persistently challenging tokens or those requiring visual grounding remain active. This dynamic gating enables the model to allocate its computational budget to the most informative parts of the sequence at each epoch, rather than committing to a fixed sparsity pattern. Finally, the per-sample binary masks are concatenated and padded to form a batch mask where is the expanded sequence length accounting for visual tokens.
3.4 Adaptive Decoder Sparsity
To exploit the binary mask during forward propagation, we modify the transformer decoder layer of the backbone LLM. We implement sparse attention by passing the mask directly as the attention mask to flash attention routines and by zeroing out the hidden states of pruned tokens. For the feed-forward network, we gather only the active positions, apply the MLP to them, and scatter the outputs back to their positions. The residual connections ensure that skipped tokens retain their previous representations. Algorithm 1 presents the pseudocode for a single forward decoder layer. This implementation requires no additional parameters and integrates seamlessly into popular libraries, such as HuggingFace Transformers. Importantly, our modifications do not affect the model architecture and thus remain compatible with pre-trained weights.
4 Experiments
4.1 Implementation Details
To demonstrate the effectiveness of the proposed framework, we apply ReGATE to two different models (i.e., VideoLLaMA2 and VideoChat2) and training strategies. We select VideoChat2 and VideoLLaMA2 over newer models like Qwen-2.5-VL and VideoLLaMA3 because ReGATE assumes access to pretrained model weights for fine-tuning. However, in many cases, these weights are not publicly available, making it infeasible to apply methods like ReGATE directly. Training such models from scratch is also impractical, as many recent MLLMs rely on proprietary pretraining pipelines that require hundreds of GPUs, web-scale datasets, and access to private data. Nonetheless, with sufficient resources and access to pretrained weights and training data, ReGATE can be seamlessly integrated into the training pipeline of any modern MLLM.
VideoLLaMA2.
We apply ReGATE to VideoLLaMA2-7B Cheng et al. (2024), whose language backbone is Qwen2-7B Yang et al. (2024). The model is initially pretrained with a frozen language backbone and subsequently fine-tuned on multimodal data. We introduce token gating during this fine-tuning stage, as the language backbone becomes trainable and can thus benefit from selective token updates. Specifically, the reference teacher model is obtained by removing the visual encoder and adapter from the VideoLLaMA2 backbone, resulting in a pure text-based LLM incapable of processing visual inputs. This teacher then computes token-wise losses, where all visual tokens have been replaced by padding.
VideoChat2.
To assess ReGATE’s effectiveness in parameter-efficient fine-tuning (PEFT) scenarios, we integrate our method into the LoRA-based Stage 3 training of VideoChat2-7B Li et al. (2024c), which uses a Mistral-7B backbone. Our key adaptation here is to make the LoRA update process itself token-selective. In a conventional setup, the loss used to update the LoRA adapters is aggregated over all tokens. In our approach, the gradients for the LoRA parameters are computed exclusively from the subset of high-importance tokens identified by ReGATE. This ensures that the parameter-efficient updates are concentrated on the most informative signals, while the original language backbone weights remain frozen. The reference teacher is derived from the text-only Mistral-7B Jiang et al. (2023) backbone, following the same procedure as described previously.
Datasets and sparsity schedule.
We fine-tune VideoLLaMA2 with and without ReGATE on the VideoChatGPT dataset Maaz et al. (2024), which is a subset of VideoLLaMA2’s official fine-tuning dataset containing approximately 300,000 instruction-response pairs. For VideoChat2, we similarly use a subset of its official fine-tuning data comprising around 2.6 million instruction pairs. Training follows the dual-cycle sparsity schedule described in Section 3.2, with parameters set to , , and . To ensure stable training at the start, we prepend a global warm-up phase of 100 iterations, during which all tokens are retained. The main hyperparameters for ReGATE include an exponential moving average (EMA) decay of and a teacher loss weighting coefficient of . All experiments are run on 4 H100 GPUs using mixed-precision training.
| Model | LLM | Tokens | ScienceQA | MME | VizWiz | POPE | |
| Open-source Models | |||||||
| InstructBLIP Dai et al. (2023) | Vicuna-7B | – | 60.5 | 254.3/1137.1 | 34.5 | 86.1 | 46.4 |
| LLaVA-1.5 Liu et al. (2024a) | Vicuna-7B | – | 66.8 | 302.1/1506.2 | 50.0 | 85.9 | 66.1 |
| Qwen-VL-Chat Bai et al. (2023) | Qwen-7B | – | 68.2 | 392.1/1467.8 | 38.9 | 74.9 | 58.2 |
| LLaVA-1.6 Liu et al. (2024a) | Vicuna-7B | – | 70.1 | – | 57.6 | 86.5 | 70.2 |
| VILA1.5 Lin et al. (2024b) | Llama-2-13B | – | 79.1 | 288.9/1429.3 | 60.6 | 84.2 | 62.8 |
| LLaVA-Next Liu et al. (2024b) | Mistral-7B | – | 73.0 | 308.9/1512.3 | – | 87.3 | 72.4 |
| LLaVA-OneVision Li et al. (2024a) | Qwen2-7B | – | 95.4 | 415.7/1577.8 | 53.0 | 87.4 | 75.4 |
| Qwen2.5-VL Bai et al. (2025) | Qwen2.5-7B | – | 89.0 | 613.9/1698.1 | – | 85.9 | 77.0 |
| Proprietary Models | |||||||
| Claude3.7-Sonnet Anthropic (2025) | – | – | 90.9 | 649.6/1189.7 | – | 82.4 | 74.3 |
| Gemini-1.5-Flash Gemini et al. (2024) | – | – | 83.3 | 488.6/1589.3 | – | 88.5 | 75.0 |
| Gemini-1.5-Pro Gemini et al. (2024) | – | – | 85.7 | 548.2/1562.4 | – | 88.2 | 76.0 |
| GPT-4o Hurst et al. (2024) | – | – | 90.1 | 719.3/1609.4 | – | 85.0 | 76.4 |
| GPT-4.1 Hurst et al. (2024) | – | – | 92.8 | 673.9/1663.6 | – | 86.4 | 78.0 |
| Models w/wo ReGATE | |||||||
| VideoChat2 | Mistral-7B | 3.93B | 40.8 | 314.6/1244.0 | 28.5 | 86.2 | 45.9 |
| VideoChat2-ReGATE | Mistral-7B | 2.22B (↓ 43.51%) | 46.6 | 360.7/1287.8 | 32.5 | 85.1 | 47.2 |
| VideoLLaMA2 | Qwen2-7B | 83.82M | 61.4 | 376.4/1474.0 | 46.8 | 86.7 | 70.4 |
| VideoLLaMA2-ReGATE | Qwen2-7B | 49.27M (↓ 41.22%) | 80.5 | 391.1/1507.1 | 48.0 | 87.5 | 70.0 |
4.2 Evaluation Benchmarks
To evaluate ReGATE, we use a diverse suite of benchmarks across image, long-video, and short-video domains. All evaluations are conducted under LMMs-Eval’s111https://github.com/EvolvingLMMs-Lab/lmms-eval settings. All benchmarks used in our evaluation follow their respective licenses and are consistent with their intended use. Below, we briefly summarize the key characteristics of each benchmark.
Image understanding.
ScienceQA Lu et al. (2022) is a multimodal science exam with 21,208 multiple-choice questions and accompanying lectures and explanations; MME Fu et al. (2024) measures perception and cognition across 14 subtasks using manually created question–answer pairs; VizWiz Gurari et al. (2018) collects real photos taken by blind users and asks questions about them and whether they are answerable; POPE Li et al. (2023) is an object hallucination benchmark formulated as a binary-choice task; and SEED-Bench Li et al. (2024b) includes 19 thousands multiple-choice questions covering both image and video modalities across 12 dimensions.
Long-video understanding.
Video-MME Fu et al. (2025) spans six primary domains and 30 subfields with videos ranging from 11 seconds to 1 hour; it integrates frames, subtitles and audio and provides 2,700 expert-annotated question–answer pairs for holistic evaluation. LongVideoBench Wu et al. (2024) contains 3,763 videos (up to an hour) and 6,678 multiple-choice questions, many of which require referring to specific temporal segments before reasoning. MLVU Zhou et al. (2025) collects long videos from diverse genres, including movies, surveillance, and egocentric recordings, and offers multiple tasks. Studies show that existing models degrade with longer context. EgoSchema Mangalam et al. (2023) comprises more than 5,000 three-minute clips from 250 hours of egocentric data, with questions requiring reasoning over much longer temporal windows than previous datasets, and current models perform far below human level.
Short-video understanding.
MVBench Li et al. (2024c) converts 20 static image tasks into dynamic video tasks, producing multiple-choice questions that probe temporal understanding. Perception Test Pătrăucean et al. (2023) consists of 11,600 real-world videos averaging 23 seconds; it evaluates perception and reasoning across six annotation types and emphasises skills such as memory, abstraction, and physics. Vinoground Zhang et al. (2024a) comprises 1,000 short video–caption pairs designed for counterfactual temporal reasoning, where even large proprietary models struggle to distinguish subtle action differences. NExT-QA Xiao et al. (2021) offers 5,440 videos and about 52,000 questions targeting causal and temporal action reasoning.
| Model | LLM | Frames | Tokens | LongVideoBench | MLVU | EgoSchema | |
| Open-source Models | |||||||
| Video-LLaVA Lin et al. (2024a) | Vicuna-7B | 8 | – | 39.9 | 39.1 | 47.3 | 38.4 |
| LLaMA-VID Li et al. (2024d) | Llama-2-7B | 1fps | – | 25.9 | – | 33.2 | 38.5 |
| LLaVA-NeXT-Video Zhang et al. (2024b) | Vicuna-7B | 32 | – | – | 43.5 | – | 43.9 |
| LLaVA-NeXT-Video Zhang et al. (2024b) | Qwen2-32B | 32 | – | 60.2 | – | 65.5 | 60.9 |
| VILA1.5 Lin et al. (2024b) | Llama-2-40B | 8 | – | 60.1 | – | 56.7 | 58.0 |
| LLaVA-OneVision Li et al. (2024a) | Qwen2-7B | 32 | – | 58.2 | 56.4 | 64.7 | 60.1 |
| Qwen2.5-VL Bai et al. (2025) | Qwen2.5-7B | – | – | 65.1 | 56.0 | 70.2 | 65.0 |
| VideoLLaMA3 Zhang et al. (2025) | Qwen2.5-7B | 1fps | – | 66.2 | 59.8 | 73.0 | 63.3 |
| Proprietary Models | |||||||
| Gemini-1.5-Flash Gemini et al. (2024) | – | – | – | 70.3 | 61.6 | – | 65.7 |
| Gemini-1.5-Pro Gemini et al. (2024) | – | – | – | 75.0 | 64.0 | – | 71.2 |
| GPT-4o Hurst et al. (2024) | – | – | – | 71.9 | 66.7 | 64.6 | 72.2 |
| Models w/wo ReGATE | |||||||
| VideoChat2 | Mistral-7B | 16 | 3.93B | 26.0 | 21.8 | 36.0 | 55.6 |
| VideoChat2-ReGATE | Mistral-7B | 16 | 2.22B (↓ 43.51%) | 32.7 | 24.3 | 40.5 | 54.8 |
| VideoLLaMA2 | Qwen2-7B | 16 | 83.82M | 53.7 | 47.7 | 53.2 | 58.2 |
| VideoLLaMA2-ReGATE | Qwen2-7B | 16 | 49.27M (↓ 41.22%) | 54.5 | 47.6 | 54.5 | 56.4 |
| Model | LLM | Frames | Tokens | MVBench | Perception | NeXT-QA | |
| Open-source Models | |||||||
| Video-LLaVA Lin et al. (2024a) | Vicuna-7B | 8 | – | 41.0 | 44.3 | 25.8 | – |
| LLaMA-VID Li et al. (2024d) | Llama-2-7B | 1fps | – | 41.9 | 44.6 | – | – |
| LLaVA-NeXT-Video Zhang et al. (2024b) | Vicuna-7B | 32 | – | 46.5 | 48.8 | 25.6 | – |
| LLaVA-NeXT-Video Zhang et al. (2024b) | Qwen2-32B | 32 | – | – | 59.4 | – | 77.3 |
| VILA1.5 Lin et al. (2024b) | Llama-2-40B | 8 | – | – | 54.0 | – | 67.9 |
| LLaVA-OneVision Li et al. (2024a) | Qwen2-7B | 32 | – | 56.7 | 57.1 | 29.4 | 79.4 |
| Qwen2.5-VL Bai et al. (2025) | Qwen2.5-7B | – | – | 69.6 | 70.5 | – | – |
| VideoLLaMA3 Zhang et al. (2025) | Qwen2.5-7B | 1fps | – | 69.7 | 72.8 | – | 84.5 |
| Proprietary Models | |||||||
| Gemini-1.5-Pro Gemini et al. (2024) | – | – | – | 60.5 | – | 22.6 | – |
| GPT-4o Hurst et al. (2024) | – | – | – | 64.6 | – | 38.2 | – |
| Models w/wo ReGATE | |||||||
| VideoChat2 | Mistral-7B | 16 | 3.93B | 55.7 | 48.4 | 22.0 | 75.2 |
| VideoChat2-ReGATE | Mistral-7B | 16 | 2.22B (↓ 43.51%) | 56.6 | 50.0 | 22.8 | 75.5 |
| VideoLLaMA2 | Qwen2-7B | 16 | 83.82M | 52.0 | 53.0 | 24.6 | 70.8 |
| VideoLLaMA2-ReGATE | Qwen2-7B | 16 | 49.27M (↓ 41.22%) | 53.6 | 54.1 | 25.2 | 70.0 |
4.3 Baseline Models
We evaluate ReGATE against a comprehensive set of baselines, including the adapted VideoLLaMA2 Cheng et al. (2024) and VideoChat2 Li et al. (2024c) models. Our comparison covers a broad range of state-of-the-art open-source models, primarily drawn from high-performing families such as LLaVA and Qwen. We also report results from proprietary models in the Google Gemini, OpenAI GPT, and Anthropic Claude series. This diverse set of baselines spans multiple LLM backbones and model sizes, ensuring a robust and meaningful comparison. The specific models evaluated across image and video tasks are listed in Tables 1, 2, and 3.
4.4 Results
Learning better: ReGATE’s accuracy gains across image and video benchmarks.
The comprehensive results presented in Tables 1, 2, and 3 show how VideoLLaMA2 and VideoChat2 perform, with and without ReGATE, across a range of image, short video, and long video understanding benchmarks. ReGATE improves performance consistently by focusing computation on the most informative tokens. For example, VideoLLaMA2-ReGATE outperforms the baseline VideoLLaMA2 on most tasks while using 41.22% fewer tokens. Similarly, VideoChat2-ReGATE achieves better results than the baseline VideoChat2 while using 43.51% fewer tokens.
On image understanding tasks that require multimodal reasoning, both models show significant gains. VideoLLaMA2-ReGATE improves by 19.1% on ScienceQA and by up to 33.1 points on MME. VideoChat2-ReGATE improves by 5.8% and 46.1 points on the same benchmarks. For long video understanding, VideoChat2-ReGATE shows strong improvements of 6.7% on VideoMME and 4.5% on MLVU. VideoLLaMA2-ReGATE also improves, though more modestly, with gains of 0.8% and 1.3% on the same tasks. Short video tasks benefit as well. VideoLLaMA2-ReGATE improves by 1.6% on MVBench and 1.1% on Perception, while VideoChat2-ReGATE gains 0.9% and 1.6%, respectively.
Overall, these results demonstrate ReGATE’s ability to adapt across diverse tasks by efficiently directing computational resources to the most important visual and semantic content.
| Model | Tokens | Train Time | Acc. (%) |
| VideoLLaMA2 | 83.82M | 32.4h | 52.0 |
| VideoLLaMA2-ReGATE | 49.27M | 26.9h | 53.6 |
| VideoLLaMA2-ReGATE | 29.32M | 16.0h | 51.9 |
| VideoChat2 | 3.93B | 37.2h | 55.7 |
| VideoChat2-ReGATE | 2.22B | 32.5h | 56.6 |
| VideoChat2-ReGATE | 1.51B | 21.6h | 55.5 |
Learning faster: ReGATE’s efficiency gains.
Table 4 presents detailed efficiency gains in token usage, training time, and accuracy on the MVBench benchmark.
For VideoLLaMA2, ReGATE closely matches the baseline accuracy (51.9% vs. 52%) in just 16.0 hours, which is less than half the time required for standard fine-tuning (32.4 hours). It does so using only 29.32 million tokens, approximately 35% of the 83.82 million tokens used by the baseline. When training is extended to 26.9 hours (still 5.5 hours less than the baseline), ReGATE processes 41.51% fewer tokens and achieves a higher accuracy of 53.6%.
For VideoChat2, which uses parameter-efficient LoRA fine-tuning, the improvements in training time are more modest. Specifically, ReGATE closely matches the baseline accuracy (55.5% vs. 55.7%) in 21.6 hours, compared to 37.2 hours for the baseline. Furthermore, when training time increases to 32.5 hours (still 4.7 hours less than the baseline), ReGATE processes 43.51% fewer tokens (2.22 billion vs. 3.93 billion) and achieves an improved accuracy of 56.6%.
This speed-up difference between VideoLLaMA2 and VideoChat2 arises from the contrast between full and LoRA fine-tuning strategies. In full fine-tuning, as used in VideoLLaMA2, both forward and backward passes through the model are computationally expensive. By pruning tokens, ReGATE speeds up both passes, especially the backward pass where gradients are computed for all model parameters. In LoRA fine-tuning, as used in VideoChat2, most parameters are frozen, and the backward pass is already efficient since gradients are only computed for a small number of adapter parameters. While ReGATE still accelerates the forward pass through the frozen backbone, the total time savings are smaller because the backward pass is not a bottleneck. Overall, ReGATE delivers significant gains in both token efficiency and training time across different training strategies, making it a flexible and effective solution for reducing computation without compromising performance.
5 Conclusion and Future Work
We introduced ReGATE, a reference-guided token gating framework that accelerates the training of multimodal large language models. By combining a student model’s learning difficulty with reference losses from a frozen text-only teacher, ReGATE dynamically focuses computation on the most informative tokens while skipping those less relevant for multimodal understanding. The method is simple to implement, requires no architectural changes, and substantially improves training efficiency. Experiments show that ReGATE achieves comparable or better accuracy than standard full fine-tuning, using only a fraction of the tokens and significantly less training time. These gains come without compromising model quality. In fact, ReGATE consistently outperforms baselines across a wide range of image and video benchmarks, demonstrating strong data efficiency and generalization. Future work will explore adaptive scheduling for token sparsity by dynamically adjusting the retained token ratio based on task complexity, model stability, and training progress (e.g., starting with higher sparsity early and relaxing it as fine-tuning progresses). We will also investigate fine-grained sparsity control across layers or attention heads for more efficient resource allocation. Moreover, we aim to generalize the notion of “reference” beyond frozen text-only teachers. Using stronger or multimodal teacher models (e.g., vision/video-language) could provide richer supervision for gating, improving cross-modal alignment and enhancing performance on complex spatial and temporal tasks.
Limitations
Due to computational resource constraints, we validate ReGATE on 7B-parameter models, VideoLLaMA2 and VideoChat2, demonstrating clear effectiveness and efficiency gains at this scale. However, the full potential of ReGATE likely emerges with larger models (e.g., 30B or 70B+ parameters) and massive, web-scale datasets, where time and cost savings become more significant. Future work should focus on evaluating ReGATE’s performance and scalability in such high-resource settings.
Appendix
Appendix A Ablation Study
How does each signal in our scoring mechanism affect performance?
To validate the contributions of the individual components within our dual-signal token scoring mechanism, we conduct an ablation study on the hyperparameter . This coefficient balances the two core signals in our difficulty score formulation: , where is the student’s dynamic EMA difficulty and is the static reference loss from the teacher model. By varying , we can isolate the impact of each signal.
We evaluate three values for : 0.0, 0.5, and 1.0. The experiments use our VideoLLaMA2-ReGATE setup with all other hyperparameters fixed for a fair comparison. As shown in Table 5, , which balances the reference loss and the student’s EMA-based difficulty, results in the best performance.
| Description | Acc. (%) | |
| Student EMA Only | 51.3 | |
| Reference Loss Only | 51.1 | |
| Combined Signals | 53.6 |
Appendix B Qualitative Analysis of Reference Loss
To validate the core mechanism of ReGATE, we qualitatively analyze the reference loss signal that guides its token selection. We assume that a high loss score from the text-only teacher indicates that a token requires visual information to be understood. Figure 3 shows two video Q&A examples, visualizing the loss for each word in the answer as calculated by a Mistral-7B Jiang et al. (2023) teacher model.
The results strongly support our assumption. As illustrated in the figure, tokens for visual details that are hard to guess from text alone, like the action “mixing” or the attribute “reflective”, get high loss scores. In contrast, simple grammatical words like “The” and “is”, or terms repeated from the question like “bartender”, get low scores. This difference confirms that reference loss is a reliable indicator of visual importance, enabling ReGATE to focus its computation on the most critical tokens for more efficient training.
| Benchmark | Response formatting prompts |
| POPE | – |
| MME | Answer the question using a single word or phrase. |
| VisWiz | Answer the question using a single word or phrase. When the provided information is insufficient, respond with “Unanswerable”. |
| ScienceQA | Answer with the option’s letter from the given choices directly. |
| SEED-Bench | Answer with the option’s letter from the given choices directly. |
| MLVU | – |
| MVBench | Only give the best option. |
| VideoMME | Answer with the option’s letter from the given choices directly. |
| EgoSchema | Answer with the option’s letter from the given choices directly. |
| NeXT-QA | – |
| Perception | Answer with the option’s letter from the given choices directly. |
| Vinoground | Please only output one English character. |
| LongVideoBench | Answer with the option’s letter from the given choices directly. |
Appendix C Additional Benchmarks Details
Table 6 lists the evaluation prompts corresponding to each benchmark used in the experiments.

### v2

@@
 Abstract
-The computational cost of training multimodal large language models (MLLMs) rapidly increases with the number of tokens involved.
-Existing efficiency methods primarily target inference and rely on token reduction or merging, offering limited benefit during training.
-In this paper, we propose ReGATE (Reference-Guided Adaptive Token Elision), an adaptive token pruning method for accelerating MLLM training.
-Specifically, ReGATE adopts a teacher-student framework in which the MLLM being trained serves as the student, and a frozen reference large language model (LLM) acts as the teacher.
-The teacher computes per-token reference losses, which are combined with an exponential moving average (EMA) of the student’s own difficulty scores.
-This adaptive difficulty-based scoring enables the selective processing of crucial tokens while bypassing less informative ones in the forward pass, significantly reducing computational overhead.
-Experiments demonstrate that ReGATE, when applied to VideoLLaMA2, matches the peak accuracy of standard training on MVBench up to 2 faster, using only 35% of the tokens.
-With additional training, it even surpasses the baseline on several multimodal benchmarks, all while reducing the total token count by over 41%.
-Code and models will be released soon.
-ReGATE: Learning Faster and Better with Fewer Tokens in MLLMs
-Chaoyu Li, Yogesh Kulkarni, Pooyan Fazli Arizona State University, Arizona, USA {chaoyuli, ykulka10, pooyan}@asu.edu
+The computational cost of training multimodal large language models (MLLMs) grows rapidly with the number of processed tokens.
+Existing efficiency methods mainly target inference via token reduction or merging, offering limited benefits during training.
+We introduce ReGATE (Reference-Guided Adaptive Token Elision), an adaptive token pruning method for accelerating MLLM training.
+ReGATE adopts a teacher-student framework, in which a frozen teacher LLM provides per-token guidance losses that are fused with an exponential moving average of the student’s difficulty estimates.
+This adaptive scoring mechanism dynamically selects informative tokens while skipping redundant ones in the forward pass, substantially reducing computation without altering the model architecture.
+Across three representative MLLMs, ReGATE matches the peak accuracy of standard training on MVBench up to 2 faster, using only 38% of the tokens.
+With extended training, it even surpasses the baseline across multiple multimodal benchmarks, cutting total token usage by over 41%.
+Code and models will be released publicly.
 1 Introduction
-Multimodal large language models (MLLMs) face significant challenges due to the high computational cost of training.
-A key bottleneck is the self-attention mechanism, whose complexity grows quadratically with input sequence length Vaswani et al.
-(2017).
-This problem is amplified in video tasks, where frames are tokenized into extremely long sequences.
-Consequently, training MLLMs on large-scale instructional datasets demands substantial computing resources, limiting accessibility and slowing progress in the field.
-Several strategies have been proposed to speed up inference in MLLMs, including static token pruning Arif et al.
-(2025) and token merging Chen et al.
-(2024).
-However, reducing the high cost of training remains a more complex and less explored challenge.
-In the unimodal text domain, recent work such as RHO-1 has introduced learnable token pruning techniques that improve training efficiency Lin et al.
-(2024c).
-Yet, these training-time acceleration methods have not been extended to large multimodal models.
-Earlier attempts to improve visual processing efficiency, typically targeting standard vision transformers (Akbari et al., 2021) or early video-language models (Lei et al., 2021), have relied on heuristic approaches such as random token dropping.
-These methods fall short in modern MLLMs, as they fail to capture the subtle and often unintuitive cross-modal importance of tokens, particularly in video, where information is both dense and temporally distributed.
-As a result, such methods risk discarding important visual or semantic content, which can lead to unstable training and weaker multimodal understanding.
-To address this challenge, we introduce ReGATE (Reference-Guided Adaptive Token Elision), a framework designed to accelerate the training of MLLMs.
-ReGATE adopts a teacher-student architecture, where the student is the multimodal model being trained, and the teacher is a frozen, text-only version of the same LLM backbone.
-This setup enables ReGATE to dynamically identify and retain the most informative tokens during training by combining two complementary signals.
-First, it assesses whether a token requires visual grounding by checking if the text-only teacher can accurately predict it from the prompt alone.
+Multimodal large language models (MLLMs) face significant challenges due to the high computational cost of training [jin2024efficient].
+A key bottleneck is the self-attention mechanism, whose complexity grows quadratically with input sequence length [nips2017_attention].
+This issue is especially severe in video tasks [videosavi, videopasta, avatar], where frames are tokenized into extremely long sequences.
+Consequently, training MLLMs on large-scale datasets requires substantial computing resources, limiting accessibility and slowing progress.
+Improving training efficiency is therefore essential for enabling MLLMs to scale to longer contexts, support richer modalities, and learn from larger datasets that would otherwise be infeasible to process.
+Several strategies have been proposed to speed up inference in MLLMs, including token pruning [ye2024atpllava, arif2025hired, alvar2025divpruned, lee2025tamp], token merging [chen2024fastv, dhouib2025pact], and token compression [yang2025libra].
+While these methods reduce FLOPs during inference, they rely on pre-scored or frozen tokens and cannot adapt token importance dynamically during training.
+As a result, reducing training costs remains a more complex and less explored problem.
+In unimodal text models, learnable token pruning methods such as RHO-1 [lin2024rho1] have improved training efficiency, but these approaches have not been extended to large multimodal models.
+Early efforts to speed up visual processing, often focusing on standard vision transformers [akbari2021vatt] or early video-language models [lei2021clipbert], rely on heuristics like random token dropping.
+In multimodal settings, however, determining which tokens are important often depends on subtle visual evidence, temporal cues, or interactions between multiple modalities.
+Ignoring these dependencies can lead to the removal of crucial information, resulting in inefficient training and reduced multimodal understanding.
+This gap highlights the need for a principled method to estimate token importance during training, one that captures cross-modal dependencies without relying on heuristics.
+To address this challenge, we introduce ReGATE (Reference‑Guided Adaptive Token Elision), a framework designed to accelerate the training of MLLMs (Figure 1).
+ReGATE adopts a teacher-student architecture, in which the student is the multimodal model being trained, and the teacher is a frozen, text-only version of the same LLM backbone.
+ReGATE combines two complementary signals to identify and retain the most informative tokens during training dynamically.
+First, it checks whether a token requires visual grounding by seeing if the text-only teacher can predict it from the prompt alone.
 Second, it evaluates the student model’s learning progress using an exponential moving average (EMA) of token-wise historical losses.
 By integrating these signals, ReGATE allocates computation to the subset of tokens that are both critical for multimodal understanding and remain challenging for the model to learn.
+To validate the effectiveness and generality of ReGATE, we evaluate it on three representative MLLMs: VideoChat2 [li2024mvbench], VideoLLaMA2 [cheng2024videollama2], and InternVL3.5 [wang2025internvl35], covering a wide range of architectures, training paradigms, and scales, and assess performance across diverse image and video benchmarks.
 To summarize, our contributions are threefold:
@@
 We show that the model-agnostic ReGATE integrates seamlessly into existing MLLMs, requiring no architectural changes, making it easy to adopt.
-Extensive experiments on image and video benchmarks demonstrate ReGATE’s broad applicability and efficiency.
-Notably, on the challenging MVBench benchmark, ReGATE, when applied to VideoLLaMA2, matches the baseline’s peak accuracy in just 16.0 hours (compared to 32.4 hours for standard fine-tuning) while processing only 29.3 million tokens, a 65% reduction from the baseline’s 83.8 million (Figure 1).
+Extensive experiments on five image benchmarks and eight video benchmarks demonstrate ReGATE’s broad applicability and efficiency.
+On the challenging MVBench benchmark, ReGATE reaches the baseline’s peak accuracy up to 2 faster while using only 38% of the tokens on average.
 2 Related Work
 2.1 Token Compression for Fast Inference
-Most existing work in the literature focus on accelerating inference, not training.
-Inference-time sparsity methods have shown that many tokens can be removed or merged with minimal impact on accuracy.
-In vision transformers, Dynamic Token PruningTang et al.
-(2023) halts processing of “easy” tokens layer by layer, reducing FLOPs by 20–35% on semantic segmentation tasks without degrading performance.
-For video LLMs, DyCokeTao et al.
-(2025) dynamically compresses spatial-temporal tokens during inference, achieving up to 2 speed-ups while keeping model weights frozen.
-Moving from pruning to aggregation, Importance-Based Token Merging Wu et al.
-(2025) merges highly similar tokens rather than dropping them, maintaining performance on long-video benchmarks while delivering 1.5 faster inference.
+Most existing work focuses on accelerating inference, not training.
+Inference-time sparsity methods show that many tokens can be removed or merged with minimal impact on accuracy [shi2025staticdynamic, jiang2025visa, yang2025topv, hyun2025multigranular].
+In vision transformers, Dynamic Token Pruning [tang2023dtop] halts processing of easy tokens layer by layer, reducing FLOPs by 20–35% without degrading performance.
+For video LLMs, DyCoke [tao2025dycoke] compresses spatial-temporal tokens during inference, achieving up to 2 speed-ups while keeping model weights frozen.
+Moving beyond pruning, Importance-Based Token Merging [wu2025imagebased] merges similar tokens rather than dropping them, maintaining performance on long-video benchmarks while delivering 1.5 faster inference.
 However, all these methods operate after training is complete.
-During training, the full token sequence is still processed in every forward and backward pass, leaving the computational cost of training mainly unaddressed.
+During training, the full token is still processed in every forward and backward pass, leaving the computational cost of training mainly unaddressed.
 2.2 Token Compression for Fast Training
 Only a few studies have explored token compression during training, rather than just at inference.
-In text-only language models, RHO-1 Lin et al.
-(2024c) ranks tokens with a reference model and backpropagates only through the most difficult subset, reducing pre-training tokens by 50% while improving accuracy.
-For MLLMs, LaVi Yue et al.
-(2025) avoids processing long visual sequences by injecting vision-conditioned deltas—small, token-specific offsets derived from the visual input—into layer norms, eliminating most visual tokens, but this requires a specialized modulation pathway that must be trained from scratch.
-LLaVA-Meteor li et al.
-(2025) introduces a flash-fusion module and a dual-expert scorer that prunes 75–95% visual tokens during instruction tuning but adds extra parameters and targets only vision tokens.
-In contrast, ReGATE uniquely combines two complementary difficulty signals: a static, cross-modal reference loss from a frozen text-only teacher that identifies tokens requiring visual grounding, and a dynamic learning signal based on the student model’s own token-wise loss tracked via an exponential moving average (EMA).
-This fusion of global and local difficulty enables a highly adaptive, parameter-free sparsity mechanism that gates both text and vision tokens, without modifying the underlying model architecture.
-2.3 Teacher-Student Distillation for MLLMs
-Most distillation approaches for MLLMs mainly focus on parameter compression.
-A systematic study Xu et al.
-(2024) shows that jointly aligning tokens and logits helps a smaller student model inherit visual grounding from a larger teacher model.
-Similarly, methods like DIME-FM Sun et al.
-(2023) show how cross-modal features can be transferred even from unpaired data.
-A more recent approach, MaskedKD Son et al.
-(2024), improves efficiency by masking a portion of the image patch tokens fed to the teacher based on the student’s attention scores.
-This strategy saves up to 50% of the teacher’s FLOPs without reducing student accuracy.
-However, MaskedKD only sparsifies the teacher’s computation and still requires backpropagation through all student tokens.
-In contrast, ReGATE introduces a fundamentally different approach by redefining the teacher’s role in distillation.
-It uses the teacher’s per-token loss to decide which tokens the student should process during each forward and backward pass.
-Instead of focusing on compressing the model itself, ReGATE targets compressing the computation path.
-This novel paradigm provides on-the-fly, modality-agnostic sparsity that optimizes the training process without changing the student’s backbone architecture.
+In text‑only language models, RHO‑1 [lin2024rho1] ranks tokens with a reference model and backpropagates only through the most difficult subset, reducing pre‑training tokens by 50% while improving accuracy.
+For MLLMs, LaVi [yue2025lavi] injects vision-conditioned deltas into layer norms to skip visual tokens but needs extra modulation layers.
+LLaVA‑Meteor [li2025llavameteror] introduces a flash‑fusion module and a dual‑expert scorer that prunes 75–95% visual tokens during instruction tuning but adds extra parameters and targets only visual tokens.
+In contrast, ReGATE combines a static cross-modal reference loss from a text-only teacher with a dynamic EMA-based student signal.
+Together, these signals produce an adaptive, parameter-free sparsity mechanism that selectively gates textual tokens while preserving visual ones, without modifying the model architecture.
+2.3 Teacher‑Student Distillation for MLLMs
+Most distillation approaches for MLLMs mainly focus on parameter compression [li2023unlock, cai2025llavakd, udandarao2025activedatacuratione].
+A systematic study [xu2024llavadi] shows that jointly aligning tokens and logits helps a smaller student model inherit visual grounding from a larger teacher model.
+Similarly, methods like DIME‑FM [sun2023dimefm] show how cross-modal features can be transferred even from unpaired data.
+More recently, MaskedKD [son2024rolemasking] reduces teacher FLOPs by up to 50% by masking patch tokens based on student attention, but it still backpropagates through all student tokens.
+In contrast, ReGATE leverages the teacher’s per-token loss to decide which tokens the student should process during each forward and backward pass.
+This approach significantly reduces computation without altering parameters and enables model-agnostic sparsity for more efficient training.
 3 ReGATE
@@
 This results in a pure text-only LLM that acts as a fixed expert to estimate the degree to which each token depends on visual input.
-Given a batch of input sequences containing both text and visual tokens, we generate a binary mask that determines which token positions should be actively computed and which can be skipped.
+Given a batch of input sequences containing both text and visual tokens, we generate a binary mask that determines which tokens should be actively computed and which can be skipped.
 This section explains how we calculate per-token difficulty scores using the frozen text-only teacher combined with the student’s own training history, how we dynamically adjust the fraction of tokens retained during training, and how we apply the resulting mask within the transformer decoder.
@@
 A higher value of indicates that token in sample has consistently posed difficulties during training.
-We then combine the reference loss and the student’s historical difficulty into a unified difficulty score for each token:
+We then combine the reference loss and the student’s historical difficulty into a unified score for each token:
 where balances these two signals.
@@
 Note that this combined difficulty evaluation is performed exclusively on output tokens (labels), as these tokens directly influence the training process through backpropagation.
-3.2 Dual-cycle Sparsity Schedule
+3.2 Dual‑cycle Sparsity Schedule
 We employ a deterministic schedule to determine the fraction of tokens kept at each training step.
@@
 This dynamic gating enables the model to allocate its computational budget to the most informative parts of the sequence at each epoch, rather than committing to a fixed sparsity pattern.
-Finally, the per-sample binary masks are concatenated and padded to form a batch mask where is the expanded sequence length accounting for visual tokens.
+Finally, the per‑sample binary masks are concatenated and padded to form a batch mask where is the expanded sequence length accounting for visual tokens.
 3.4 Adaptive Decoder Sparsity
@@
 We implement sparse attention by passing the mask directly as the attention mask to flash attention routines and by zeroing out the hidden states of pruned tokens.
-For the feed-forward network, we gather only the active positions, apply the MLP to them, and scatter the outputs back to their positions.
+For the feed‑forward network, we gather only the active positions, apply the MLP to them, and scatter the outputs back to their positions.
 The residual connections ensure that skipped tokens retain their previous representations.
-Algorithm 1 presents the pseudocode for a single forward decoder layer.
-This implementation requires no additional parameters and integrates seamlessly into popular libraries, such as HuggingFace Transformers.
-Importantly, our modifications do not affect the model architecture and thus remain compatible with pre-trained weights.
+A single decoder layer forward pass applies self-attention and MLP only to tokens selected by the binary mask, while skipped tokens are propagated unchanged via residual connections.
+This design introduces no additional parameters, integrates seamlessly with libraries such as HuggingFace Transformers, and remains fully compatible with pre-trained weights.
+During back propagation, gradients are computed only for parameters associated with the kept tokens.
+Skipped tokens are treated as constants.
+Their activations and losses receive no gradient updates, while residual connections ensure that gradient flow remains continuous across layers.
+This masking rule aligns the backward graph with the sparse forward path and maintains training stability in practice.
+Complete pseudocode and implementation details (including Q/K/V masking) are provided in Supplementary Material.
 4 Experiments
+| Model | LLM | Tokens | ScienceQA | MME | VizWiz | POPE | |
+[…]
+| InstructBLIP [dai2023instructblip] | Vicuna-7B | – | 60.5 | 254.3/1137.1 | 34.5 | 86.1 | 46.4 |
+| LLaVA-1.5 [liu2024llava] | Vicuna-7B | – | 66.8 | 302.1/1506.2 | 50.0 | 85.9 | 66.1 |
+| Qwen-VL-Chat [bai2023qwenvl] | Qwen-7B | – | 68.2 | 392.1/1467.8 | 38.9 | 74.9 | 58.2 |
+| LLaVA-1.6 [liu2024llava] | Vicuna-7B | – | 70.1 | – | 57.6 | 86.5 | 70.2 |
+| VILA1.5 [lin2024vila] | Llama-2-13B | – | 79.1 | 288.9/1429.3 | 60.6 | 84.2 | 62.8 |
+| LLaVA-Next [liu2024llavanext] | Mistral-7B | – | 73.0 | 308.9/1512.3 | – | 87.3 | 72.4 |
+| LLaVA-OneVision [llavaov] | Qwen2-7B | – | 95.4 | 415.7/1577.8 | 53.0 | 87.4 | 75.4 |
+| Qwen2.5-VL [bai2025qwen25vl] | Qwen2.5-7B | – | 89.0 | 613.9/1698.1 | – | 85.9 | 77.0 |
+[…]
+| Claude3.7-Sonnet [anthropic2025claude37] | – | – | 90.9 | 649.6/1189.7 | – | 82.4 | 74.3 |
+| Gemini-1.5-Flash [geminiteam2024gemini15] | – | – | 83.3 | 488.6/1589.3 | – | 88.5 | 75.0 |
+| Gemini-1.5-Pro [geminiteam2024gemini15] | – | – | 85.7 | 548.2/1562.4 | – | 88.2 | 76.0 |
+| GPT-4o [openai2024gpt4ocard] | – | – | 90.1 | 719.3/1609.4 | – | 85.0 | 76.4 |
+| GPT-4.1 [openai2024gpt4ocard] | – | – | 92.8 | 673.9/1663.6 | – | 86.4 | 78.0 |
+[…]
+| VideoChat2 | Mistral-7B | 3.93B | 40.8 | 314.6/1244.0 | 28.5 | 86.2 | 45.9 |
+[…]
+| VideoLLaMA2 | Qwen2-7B | 83.82M | 61.4 | 376.4/1474.0 | 46.8 | 86.7 | 70.4 |
+[…]
+| InternVL3.5 | Qwen3-14B | 3.96B | 93.3 | 681.6/1694.3 | 60.6 | 91.6 | 76.8 |
+[…]
 4.1 Implementation Details
-To demonstrate the effectiveness of the proposed framework, we apply ReGATE to two different models (i.e., VideoLLaMA2 and VideoChat2) and training strategies.
-We select VideoChat2 and VideoLLaMA2 over newer models like Qwen-2.5-VL and VideoLLaMA3 because ReGATE assumes access to pretrained model weights for fine-tuning.
-However, in many cases, these weights are not publicly available, making it infeasible to apply methods like ReGATE directly.
-Training such models from scratch is also impractical, as many recent MLLMs rely on proprietary pretraining pipelines that require hundreds of GPUs, web-scale datasets, and access to private data.
-Nonetheless, with sufficient resources and access to pretrained weights and training data, ReGATE can be seamlessly integrated into the training pipeline of any modern MLLM.
+To demonstrate the effectiveness and generality of ReGATE, we apply it to three representative MLLMs, VideoChat2 [li2024mvbench], VideoLLaMA2 [cheng2024videollama2], and InternVL3.5 [wang2025internvl35], spanning diverse architectures, training paradigms, and scales.
+We exclude models such as Qwen2.5-VL [bai2025qwen25vl] and VideoLLaMA3 [zhang2025videollama3] due to the lack of publicly available pretrained weights.
+Training such models from scratch is impractical, as they rely on web-scale data and hundreds of GPUs.
+However, with sufficient resources and pretrained weights, ReGATE can be seamlessly integrated into any MLLM training pipeline.
+For all experiments, the per-token reference losses from the text-only teacher are computed once over the entire fine-tuning dataset before student training begins and then cached and reused throughout training.
 VideoLLaMA2.
-We apply ReGATE to VideoLLaMA2-7B Cheng et al.
-(2024), whose language backbone is Qwen2-7B Yang et al.
-(2024).
-The model is initially pretrained with a frozen language backbone and subsequently fine-tuned on multimodal data.
-We introduce token gating during this fine-tuning stage, as the language backbone becomes trainable and can thus benefit from selective token updates.
-Specifically, the reference teacher model is obtained by removing the visual encoder and adapter from the VideoLLaMA2 backbone, resulting in a pure text-based LLM incapable of processing visual inputs.
-This teacher then computes token-wise losses, where all visual tokens have been replaced by padding.
+We first apply ReGATE to VideoLLaMA2-7B [cheng2024videollama2], whose Qwen2-7B [yang2024qwen2] backbone is unfrozen during multimodal fine-tuning.
+ReGATE is introduced at this stage, using a text-only reference teacher (the same backbone without the visual encoder or adapter) that computes per-token losses with visual tokens masked out.
+This configuration allows us to examine how ReGATE interacts with full fine-tuning on a standard multimodal backbone.
 VideoChat2.
-To assess ReGATE’s effectiveness in parameter-efficient fine-tuning (PEFT) scenarios, we integrate our method into the LoRA-based Stage 3 training of VideoChat2-7B Li et al.
-(2024c), which uses a Mistral-7B backbone.
-Our key adaptation here is to make the LoRA update process itself token-selective.
-In a conventional setup, the loss used to update the LoRA adapters is aggregated over all tokens.
-In our approach, the gradients for the LoRA parameters are computed exclusively from the subset of high-importance tokens identified by ReGATE.
-This ensures that the parameter-efficient updates are concentrated on the most informative signals, while the original language backbone weights remain frozen.
-The reference teacher is derived from the text-only Mistral-7B Jiang et al.
-(2023) backbone, following the same procedure as described previously.
+To evaluate ReGATE under parameter-efficient fine-tuning (PEFT), we integrate it into the LoRA-based Stage 3 training of VideoChat2-7B [li2024mvbench], built on Mistral-7B [jiang2023mistral7b].
+Gradients for LoRA parameters are computed only from the high-importance tokens selected by ReGATE, while the language backbone weights remain frozen.
+This setup enables us to assess whether ReGATE can complement PEFT methods without architectural modifications.
+InternVL3.5.
+Finally, we evaluate scalability of ReGATE on InternVL3.5-14B [wang2025internvl35], based on Qwen3-14B [yang2025qwen3technicalreport].
+The reference teacher is derived from the same backbone by removing its visual encoder and projector, forming a text-only LLM.
+During fine-tuning, ReGATE dynamically gates tokens to focus computation on the most informative positions, reducing activation memory and training FLOPs.
+This setup allows us to assess ReGATE’s scalability and stability at large model scales.
 Datasets and sparsity schedule.
-We fine-tune VideoLLaMA2 with and without ReGATE on the VideoChatGPT dataset Maaz et al.
-(2024), which is a subset of VideoLLaMA2’s official fine-tuning dataset containing approximately 300,000 instruction-response pairs.
+We fine-tune VideoLLaMA2 with and without ReGATE on the VideoChatGPT dataset [maaz2024videochatgpt], which is a subset of VideoLLaMA2’s official fine-tuning dataset containing approximately 300,000 instruction-response pairs.
 For VideoChat2, we similarly use a subset of its official fine-tuning data comprising around 2.6 million instruction pairs.
+For InternVL3.5, we adopt the same 2.6M-sample dataset used for VideoChat2, since the full InternVL3.5 training corpus is extremely large and not fully released.
 Training follows the dual-cycle sparsity schedule described in Section 3.2, with parameters set to , , and .
@@
 The main hyperparameters for ReGATE include an exponential moving average (EMA) decay of and a teacher loss weighting coefficient of .
-All experiments are run on 4 H100 GPUs using mixed-precision training.
-| Model | LLM | Tokens | ScienceQA | MME | VizWiz | POPE | |
-| Open-source Models | |||||||
-| InstructBLIP Dai et al. (2023) | Vicuna-7B | – | 60.5 | 254.3/1137.1 | 34.5 | 86.1 | 46.4 |
-| LLaVA-1.5 Liu et al. (2024a) | Vicuna-7B | – | 66.8 | 302.1/1506.2 | 50.0 | 85.9 | 66.1 |
-| Qwen-VL-Chat Bai et al. (2023) | Qwen-7B | – | 68.2 | 392.1/1467.8 | 38.9 | 74.9 | 58.2 |
-| LLaVA-1.6 Liu et al. (2024a) | Vicuna-7B | – | 70.1 | – | 57.6 | 86.5 | 70.2 |
-| VILA1.5 Lin et al. (2024b) | Llama-2-13B | – | 79.1 | 288.9/1429.3 | 60.6 | 84.2 | 62.8 |
-| LLaVA-Next Liu et al. (2024b) | Mistral-7B | – | 73.0 | 308.9/1512.3 | – | 87.3 | 72.4 |
-| LLaVA-OneVision Li et al. (2024a) | Qwen2-7B | – | 95.4 | 415.7/1577.8 | 53.0 | 87.4 | 75.4 |
-| Qwen2.5-VL Bai et al. (2025) | Qwen2.5-7B | – | 89.0 | 613.9/1698.1 | – | 85.9 | 77.0 |
-| Proprietary Models | |||||||
-| Claude3.7-Sonnet Anthropic (2025) | – | – | 90.9 | 649.6/1189.7 | – | 82.4 | 74.3 |
-| Gemini-1.5-Flash Gemini et al. (2024) | – | – | 83.3 | 488.6/1589.3 | – | 88.5 | 75.0 |
-| Gemini-1.5-Pro Gemini et al. (2024) | – | – | 85.7 | 548.2/1562.4 | – | 88.2 | 76.0 |
-| GPT-4o Hurst et al. (2024) | – | – | 90.1 | 719.3/1609.4 | – | 85.0 | 76.4 |
-| GPT-4.1 Hurst et al. (2024) | – | – | 92.8 | 673.9/1663.6 | – | 86.4 | 78.0 |
-| Models w/wo ReGATE | |||||||
-| VideoChat2 | Mistral-7B | 3.93B | 40.8 | 314.6/1244.0 | 28.5 | 86.2 | 45.9 |
-| VideoChat2-ReGATE | Mistral-7B | 2.22B (↓ 43.51%) | 46.6 | 360.7/1287.8 | 32.5 | 85.1 | 47.2 |
-| VideoLLaMA2 | Qwen2-7B | 83.82M | 61.4 | 376.4/1474.0 | 46.8 | 86.7 | 70.4 |
-| VideoLLaMA2-ReGATE | Qwen2-7B | 49.27M (↓ 41.22%) | 80.5 | 391.1/1507.1 | 48.0 | 87.5 | 70.0 |
-4.2 Evaluation Benchmarks
-To evaluate ReGATE, we use a diverse suite of benchmarks across image, long-video, and short-video domains.
-All evaluations are conducted under LMMs-Eval’s111https://github.com/EvolvingLMMs-Lab/lmms-eval settings.
-All benchmarks used in our evaluation follow their respective licenses and are consistent with their intended use.
-Below, we briefly summarize the key characteristics of each benchmark.
-Image understanding.
-ScienceQA Lu et al.
-(2022) is a multimodal science exam with 21,208 multiple-choice questions and accompanying lectures and explanations; MME Fu et al.
-(2024) measures perception and cognition across 14 subtasks using manually created question–answer pairs; VizWiz Gurari et al.
-(2018) collects real photos taken by blind users and asks questions about them and whether they are answerable; POPE Li et al.
-(2023) is an object hallucination benchmark formulated as a binary-choice task; and SEED-Bench Li et al.
-(2024b) includes 19 thousands multiple-choice questions covering both image and video modalities across 12 dimensions.
-Long-video understanding.
-Video-MME Fu et al.
-(2025) spans six primary domains and 30 subfields with videos ranging from 11 seconds to 1 hour; it integrates frames, subtitles and audio and provides 2,700 expert-annotated question–answer pairs for holistic evaluation.
-LongVideoBench Wu et al.
-(2024) contains 3,763 videos (up to an hour) and 6,678 multiple-choice questions, many of which require referring to specific temporal segments before reasoning.
-MLVU Zhou et al.
-(2025) collects long videos from diverse genres, including movies, surveillance, and egocentric recordings, and offers multiple tasks.
-Studies show that existing models degrade with longer context.
-EgoSchema Mangalam et al.
-(2023) comprises more than 5,000 three-minute clips from 250 hours of egocentric data, with questions requiring reasoning over much longer temporal windows than previous datasets, and current models perform far below human level.
-Short-video understanding.
-MVBench Li et al.
-(2024c) converts 20 static image tasks into dynamic video tasks, producing multiple-choice questions that probe temporal understanding.
-Perception Test Pătrăucean et al.
-(2023) consists of 11,600 real-world videos averaging 23 seconds; it evaluates perception and reasoning across six annotation types and emphasises skills such as memory, abstraction, and physics.
-Vinoground Zhang et al.
-(2024a) comprises 1,000 short video–caption pairs designed for counterfactual temporal reasoning, where even large proprietary models struggle to distinguish subtle action differences.
-NExT-QA Xiao et al.
-(2021) offers 5,440 videos and about 52,000 questions targeting causal and temporal action reasoning.
+All experiments on VideoLLaMA2 and VideoChat2 are run on 4 H100 GPUs, while experiments on InternVL3.5 use 16 H100 GPUs, all under mixed-precision training.
+4.2 Benchmarks and Baselines
+We evaluate ReGATE across image, long-video, and short-video understanding benchmarks under the LMMs-Eval [zhang2024lmmseval] framework.
+For image understanding, we include ScienceQA [lu2022scienceqa], MME [fu2024mme], VizWiz [gurari2018vizwiz], POPE [li2023pope], and SEED [li2024seed], which together probe scientific reasoning, perception, answerability, hallucination, and broad multimodal comprehension.
+For long-video understanding, we evaluate on Video-MME [fu2025videomme], LongVideoBench [wu2024longvideobench], MLVU [zhou2025mlvu], and EgoSchema [mangalam2023egoschema], all of which emphasize extended temporal reasoning, consistency, and grounding across minutes- to hour-long sequences.
+For short-video understanding, we adopt MVBench [li2024mvbench], Perception Test [pătrăucean2023perceptiontest], Vinoground [zhang2024vinoground], and NExT-QA [xiao2021nextqa], which target fine-grained event recognition, local temporal relations, counterfactual reasoning, and causal/temporal action reasoning in clips of tens of seconds.
+We evaluate ReGATE against a diverse set of state-of-the-art models, including high-performing open-source families (LLaVA, Qwen) and proprietary models (Google Gemini, OpenAI GPT, Anthropic Claude).
+This selection spans multiple LLM backbones and sizes, ensuring robust comparisons across image and video tasks (Tables 4, 4.2, LABEL:tab:short-video).
 | Model | LLM | Frames | Tokens | LongVideoBench | MLVU | EgoSchema | |
-| Open-source Models | |||||||
-| Video-LLaVA Lin et al. (2024a) | Vicuna-7B | 8 | – | 39.9 | 39.1 | 47.3 | 38.4 |
-| LLaMA-VID Li et al. (2024d) | Llama-2-7B | 1fps | – | 25.9 | – | 33.2 | 38.5 |
-| LLaVA-NeXT-Video Zhang et al. (2024b) | Vicuna-7B | 32 | – | – | 43.5 | – | 43.9 |
-| LLaVA-NeXT-Video Zhang et al. (2024b) | Qwen2-32B | 32 | – | 60.2 | – | 65.5 | 60.9 |
-| VILA1.5 Lin et al. (2024b) | Llama-2-40B | 8 | – | 60.1 | – | 56.7 | 58.0 |
-| LLaVA-OneVision Li et al. (2024a) | Qwen2-7B | 32 | – | 58.2 | 56.4 | 64.7 | 60.1 |
-| Qwen2.5-VL Bai et al. (2025) | Qwen2.5-7B | – | – | 65.1 | 56.0 | 70.2 | 65.0 |
-| VideoLLaMA3 Zhang et al. (2025) | Qwen2.5-7B | 1fps | – | 66.2 | 59.8 | 73.0 | 63.3 |
-| Proprietary Models | |||||||
-| Gemini-1.5-Flash Gemini et al. (2024) | – | – | – | 70.3 | 61.6 | – | 65.7 |
-| Gemini-1.5-Pro Gemini et al. (2024) | – | – | – | 75.0 | 64.0 | – | 71.2 |
-| GPT-4o Hurst et al. (2024) | – | – | – | 71.9 | 66.7 | 64.6 | 72.2 |
-| Models w/wo ReGATE | |||||||
+[…]
+| Video-LLaVA [lin2024videollava] | Vicuna-7B | 8 | – | 39.9 | 39.1 | 47.3 | 38.4 |
+| LLaMA-VID [li2024llamavid] | Llama-2-7B | 1fps | – | 25.9 | – | 33.2 | 38.5 |
+| LLaVA-NeXT-Video [zhang2024llavanextvideo] | Vicuna-7B | 32 | – | – | 43.5 | – | 43.9 |
+| LLaVA-NeXT-Video [zhang2024llavanextvideo] | Qwen2-32B | 32 | – | 60.2 | – | 65.5 | 60.9 |
+| VILA1.5 [lin2024vila] | Llama-2-40B | 8 | – | 60.1 | – | 56.7 | 58.0 |
+| LLaVA-OneVision [llavaov] | Qwen2-7B | 32 | – | 58.2 | 56.4 | 64.7 | 60.1 |
+| Qwen2.5-VL [bai2025qwen25vl] | Qwen2.5-7B | – | – | 65.1 | 56.0 | 70.2 | 65.0 |
+| VideoLLaMA3 [zhang2025videollama3] | Qwen2.5-7B | 1fps | – | 66.2 | 59.8 | 73.0 | 63.3 |
+[…]
+| Gemini-1.5-Flash [geminiteam2024gemini15] | – | – | – | 70.3 | 61.6 | – | 65.7 |
+| Gemini-1.5-Pro [geminiteam2024gemini15] | – | – | – | 75.0 | 64.0 | – | 71.2 |
+| GPT-4o [openai2024gpt4ocard] | – | – | – | 71.9 | 66.7 | 64.6 | 72.2 |
+[…]
 | VideoChat2 | Mistral-7B | 16 | 3.93B | 26.0 | 21.8 | 36.0 | 55.6 |
-| VideoChat2-ReGATE | Mistral-7B | 16 | 2.22B (↓ 43.51%) | 32.7 | 24.3 | 40.5 | 54.8 |
+[…]
 | VideoLLaMA2 | Qwen2-7B | 16 | 83.82M | 53.7 | 47.7 | 53.2 | 58.2 |
-| VideoLLaMA2-ReGATE | Qwen2-7B | 16 | 49.27M (↓ 41.22%) | 54.5 | 47.6 | 54.5 | 56.4 |
-| Model | LLM | Frames | Tokens | MVBench | Perception | NeXT-QA | |
-| Open-source Models | |||||||
-| Video-LLaVA Lin et al. (2024a) | Vicuna-7B | 8 | – | 41.0 | 44.3 | 25.8 | – |
-| LLaMA-VID Li et al. (2024d) | Llama-2-7B | 1fps | – | 41.9 | 44.6 | – | – |
-| LLaVA-NeXT-Video Zhang et al. (2024b) | Vicuna-7B | 32 | – | 46.5 | 48.8 | 25.6 | – |
-| LLaVA-NeXT-Video Zhang et al. (2024b) | Qwen2-32B | 32 | – | – | 59.4 | – | 77.3 |
-| VILA1.5 Lin et al. (2024b) | Llama-2-40B | 8 | – | – | 54.0 | – | 67.9 |
-| LLaVA-OneVision Li et al. (2024a) | Qwen2-7B | 32 | – | 56.7 | 57.1 | 29.4 | 79.4 |
-| Qwen2.5-VL Bai et al. (2025) | Qwen2.5-7B | – | – | 69.6 | 70.5 | – | – |
-| VideoLLaMA3 Zhang et al. (2025) | Qwen2.5-7B | 1fps | – | 69.7 | 72.8 | – | 84.5 |
-| Proprietary Models | |||||||
-| Gemini-1.5-Pro Gemini et al. (2024) | – | – | – | 60.5 | – | 22.6 | – |
-| GPT-4o Hurst et al. (2024) | – | – | – | 64.6 | – | 38.2 | – |
-| Models w/wo ReGATE | |||||||
-| VideoChat2 | Mistral-7B | 16 | 3.93B | 55.7 | 48.4 | 22.0 | 75.2 |
-| VideoChat2-ReGATE | Mistral-7B | 16 | 2.22B (↓ 43.51%) | 56.6 | 50.0 | 22.8 | 75.5 |
-| VideoLLaMA2 | Qwen2-7B | 16 | 83.82M | 52.0 | 53.0 | 24.6 | 70.8 |
-| VideoLLaMA2-ReGATE | Qwen2-7B | 16 | 49.27M (↓ 41.22%) | 53.6 | 54.1 | 25.2 | 70.0 |
-4.3 Baseline Models
-We evaluate ReGATE against a comprehensive set of baselines, including the adapted VideoLLaMA2 Cheng et al.
-(2024) and VideoChat2 Li et al.
-(2024c) models.
-Our comparison covers a broad range of state-of-the-art open-source models, primarily drawn from high-performing families such as LLaVA and Qwen.
-We also report results from proprietary models in the Google Gemini, OpenAI GPT, and Anthropic Claude series.
-This diverse set of baselines spans multiple LLM backbones and model sizes, ensuring a robust and meaningful comparison.
-The specific models evaluated across image and video tasks are listed in Tables 1, 2, and 3.
+[…]
+| InternVL3.5 | Qwen3-14B | 16 | 3.96B | 62.4 | 57.9 | 63.7 | 64.7 |
+[…]
 4.3 Results
 Learning better: ReGATE’s accuracy gains across image and video benchmarks.
-The comprehensive results presented in Tables 1, 2, and 3 show how VideoLLaMA2 and VideoChat2 perform, with and without ReGATE, across a range of image, short video, and long video understanding benchmarks.
+The comprehensive results presented in Tables 4, 4.2, and LABEL:tab:short-video show how VideoLLaMA2, VideoChat2, and InternVL3.5 perform, with and without ReGATE, across a range of image, short video, and long video understanding benchmarks.
 ReGATE improves performance consistently by focusing computation on the most informative tokens.
+Figure LABEL:fig:saliency visualizes the learned attention maps, showing that ReGATE concentrates attention on visually and semantically important tokens compared to standard fine-tuning.
 For example, VideoLLaMA2-ReGATE outperforms the baseline VideoLLaMA2 on most tasks while using 41.22% fewer tokens.
 Similarly, VideoChat2-ReGATE achieves better results than the baseline VideoChat2 while using 43.51% fewer tokens.
-On image understanding tasks that require multimodal reasoning, both models show significant gains.
+At the larger 14B scale, InternVL3.5-ReGATE also shows gains across both image and video benchmarks, confirming that ReGATE remains effective for high-capacity backbones.
+On image understanding tasks requiring multimodal reasoning, all three models show significant gains.
 VideoLLaMA2-ReGATE improves by 19.1% on ScienceQA and by up to 33.1 points on MME.
 VideoChat2-ReGATE improves by 5.8% and 46.1 points on the same benchmarks.
-For long video understanding, VideoChat2-ReGATE shows strong improvements of 6.7% on VideoMME and 4.5% on MLVU.
+InternVL3.5-ReGATE also shows clear improvements, including +7.7/+4.5 points on MME, 0.9% on VizWiz, and 1.6% on POPE.
+For long-video understanding, VideoChat2-ReGATE shows strong improvements of 6.7% on Video-MME and 4.5% on MLVU.
 VideoLLaMA2-ReGATE also improves, though more modestly, with gains of 0.8% and 1.3% on the same tasks.
+At the larger scale, InternVL3.5-ReGATE further improves performance by 0.6% on Video-MME, and 0.5% on MLVU.
 Short video tasks benefit as well.
-VideoLLaMA2-ReGATE improves by 1.6% on MVBench and 1.1% on Perception, while VideoChat2-ReGATE gains 0.9% and 1.6%, respectively.
-Overall, these results demonstrate ReGATE’s ability to adapt across diverse tasks by efficiently directing computational resources to the most important visual and semantic content.
-| Model | Tokens | Train Time | Acc. (%) |
-| VideoLLaMA2 | 83.82M | 32.4h | 52.0 |
-| VideoLLaMA2-ReGATE | 49.27M | 26.9h | 53.6 |
-| VideoLLaMA2-ReGATE | 29.32M | 16.0h | 51.9 |
-| VideoChat2 | 3.93B | 37.2h | 55.7 |
-| VideoChat2-ReGATE | 2.22B | 32.5h | 56.6 |
-| VideoChat2-ReGATE | 1.51B | 21.6h | 55.5 |
-Learning faster: ReGATE’s efficiency gains.
-Table 4 presents detailed efficiency gains in token usage, training time, and accuracy on the MVBench benchmark.
-For VideoLLaMA2, ReGATE closely matches the baseline accuracy (51.9% vs.
-52%) in just 16.0 hours, which is less than half the time required for standard fine-tuning (32.4 hours).
-It does so using only 29.32 million tokens, approximately 35% of the 83.82 million tokens used by the baseline.
-When training is extended to 26.9 hours (still 5.5 hours less than the baseline), ReGATE processes 41.51% fewer tokens and achieves a higher accuracy of 53.6%.
-For VideoChat2, which uses parameter-efficient LoRA fine-tuning, the improvements in training time are more modest.
-Specifically, ReGATE closely matches the baseline accuracy (55.5% vs.
-55.7%) in 21.6 hours, compared to 37.2 hours for the baseline.
-Furthermore, when training time increases to 32.5 hours (still 4.7 hours less than the baseline), ReGATE processes 43.51% fewer tokens (2.22 billion vs.
-3.93 billion) and achieves an improved accuracy of 56.6%.
-This speed-up difference between VideoLLaMA2 and VideoChat2 arises from the contrast between full and LoRA fine-tuning strategies.
-In full fine-tuning, as used in VideoLLaMA2, both forward and backward passes through the model are computationally expensive.
-By pruning tokens, ReGATE speeds up both passes, especially the backward pass where gradients are computed for all model parameters.
-In LoRA fine-tuning, as used in VideoChat2, most parameters are frozen, and the backward pass is already efficient since gradients are only computed for a small number of adapter parameters.
-While ReGATE still accelerates the forward pass through the frozen backbone, the total time savings are smaller because the backward pass is not a bottleneck.
-Overall, ReGATE delivers significant gains in both token efficiency and training time across different training strategies, making it a flexible and effective solution for reducing computation without compromising performance.
-5 Conclusion and Future Work
-We introduced ReGATE, a reference-guided token gating framework that accelerates the training of multimodal large language models.
-By combining a student model’s learning difficulty with reference losses from a frozen text-only teacher, ReGATE dynamically focuses computation on the most informative tokens while skipping those less relevant for multimodal understanding.
-The method is simple to implement, requires no architectural changes, and substantially improves training efficiency.
-Experiments show that ReGATE achieves comparable or better accuracy than standard full fine-tuning, using only a fraction of the tokens and significantly less training time.
-These gains come without compromising model quality.
-In fact, ReGATE consistently outperforms baselines across a wide range of image and video benchmarks, demonstrating strong data efficiency and generalization.
-Future work will explore adaptive scheduling for token sparsity by dynamically adjusting the retained token ratio based on task complexity, model stability, and training progress (e.g., starting with higher sparsity early and relaxing it as fine-tuning progresses).
-We will also investigate fine-grained sparsity control across layers or attention heads for more efficient resource allocation.
-Moreover, we aim to generalize the notion of “reference” beyond frozen text-only teachers.
-Using stronger or multimodal teacher models (e.g., vision/video-language) could provide richer supervision for gating, improving cross-modal alignment and enhancing performance on complex spatial and temporal tasks.
-Limitations
-Due to computational resource constraints, we validate ReGATE on 7B-parameter models, VideoLLaMA2 and VideoChat2, demonstrating clear effectiveness and efficiency gains at this scale.
-However, the full potential of ReGATE likely emerges with larger models (e.g., 30B or 70B+ parameters) and massive, web-scale datasets, where time and cost savings become more significant.
-Future work should focus on evaluating ReGATE’s performance and scalability in such high-resource settings.
-Appendix
-Appendix A Ablation Study
-How does each signal in our scoring mechanism affect performance?
+VideoLLaMA2-ReGATE improves by 1.6% on MVBench and 1.1% on Perception, and VideoChat2-ReGATE gains 0.9% and 1.6%.
+InternVL3.5-ReGATE also shows improvements, with +1.3% on MVBench and +1.4% on Perception.
+Beyond performance metrics, ReGATE generalizes well across architectures and training strategies.
+It works effectively with transformer backbones such as Mistral, Qwen2, and Qwen3, and retains its benefits under both full-parameter and LoRA fine-tuning.
+These results confirm that ReGATE’s efficiency and accuracy gains are architecture- and training-strategy–agnostic, enabling seamless integration into diverse multimodal models.
+Supplementary Material
+6 Sparse Decoder Forward and Masking Details
+Algorithm 1 shows how the binary mask controls computation within each decoder layer.
+The hidden states are first normalized, and self-attention is applied only to tokens marked active in the mask.
+The outputs of attention are added back through the residual connection, after which the MLP is executed solely on the active positions.
+Finally, the updated values are written back to the corresponding slots, while masked tokens remain unchanged and are propagated forward through residual paths.
+This ensures that only informative tokens are processed, while inactive ones are efficiently skipped without disrupting model integrity.
+Q/K/V Masking.
+For clarity, we specify how the binary mask is applied within transformer layers.
+During self-attention, queries (Q) are computed only for the active tokens, while keys (K) and values (V) are restricted to the same set, with pruned tokens excluded from the gather step.
+In the feed-forward network (FFN), hidden states at active positions are passed through the MLP and the outputs are written back to their original slots.
+Residual connections ensure that skipped tokens simply propagate unchanged across layers.
+This design realizes sparse computation while leaving the model parameters intact and fully compatible with standard transformer implementations.
+7 Additional Ablation Studies
+In this section, we delve deeper into the design choices of ReGATE, specifically examining the sensitivity of the weighting hyperparameter and the impact of the reference teacher’s capacity.
+7.1 Impact of the Weighting Factor
 To validate the contributions of the individual components within our dual-signal token scoring mechanism, we conduct an ablation study on the hyperparameter .
 This coefficient balances the two core signals in our difficulty score formulation: , where is the student’s dynamic EMA difficulty and is the static reference loss from the teacher model.
-By varying , we can isolate the impact of each signal.
 We evaluate three values for : 0.0, 0.5, and 1.0.
 The experiments use our VideoLLaMA2-ReGATE setup with all other hyperparameters fixed for a fair comparison.
-As shown in Table 5, , which balances the reference loss and the student’s EMA-based difficulty, results in the best performance.
+As shown in Table 7, setting achieves the best balance.
+Relying solely on the student’s historical difficulty () fails to capture zero-shot visual dependencies, while relying exclusively on the teacher () ignores the student’s training dynamics.
+The combined signal effectively isolates tokens that are both visually critical and challenging for the current model state.
 | Description | Acc. (%) | |
-| Student EMA Only | 51.3 | |
-| Reference Loss Only | 51.1 | |
-| Combined Signals | 53.6 |
-Appendix B Qualitative Analysis of Reference Loss
-To validate the core mechanism of ReGATE, we qualitatively analyze the reference loss signal that guides its token selection.
-We assume that a high loss score from the text-only teacher indicates that a token requires visual information to be understood.
-Figure 3 shows two video Q&A examples, visualizing the loss for each word in the answer as calculated by a Mistral-7B Jiang et al.
-(2023) teacher model.
-The results strongly support our assumption.
-As illustrated in the figure, tokens for visual details that are hard to guess from text alone, like the action “mixing” or the attribute “reflective”, get high loss scores.
-In contrast, simple grammatical words like “The” and “is”, or terms repeated from the question like “bartender”, get low scores.
-This difference confirms that reference loss is a reliable indicator of visual importance, enabling ReGATE to focus its computation on the most critical tokens for more efficient training.
+| Student EMA Only | 47.7 | |
+| Reference Loss Only | 46.4 | |
+| Combined Signals | 48.9 |
+7.2 Impact of Reference Teacher Capacity
+| Teacher Model | Avg. Acc. (%) |
+| Qwen2-1.5B | 45.4 |
+| Qwen2-57B | 46.8 |
+| Qwen2-7B (Ours) | 48.9 |
+A critical design choice in ReGATE is the selection of the reference teacher.
+We investigate how the capacity of the teacher model affects the student’s performance.
+Specifically, for the VideoLLaMA2-7B [cheng2024videollama2] student (based on Qwen2-7B [yang2024qwen2]), we evaluate reference signals from a smaller model (Qwen2-1.5B) and a substantially larger model (Qwen2-57B-A14B), comparing them against the matched Qwen2-7B baseline.
+To ensure fair comparison, we restrict the study to a single model family, which guarantees consistent tokenization and avoids misalignment between teacher and student.
+Cross-architecture teachers (e.g., Mistral guiding Qwen) are excluded because distinct tokenizers yield inconsistent token boundaries for the same text, violating the one-to-one mapping required for per-token reference loss computation in ReGATE.
+Such mismatch renders the reference signal mathematically invalid.
+The results in Table 8 show that the matched Qwen2-7B teacher outperforms both the smaller 1.5B and the much larger 57B models.
+This pattern reveals a capacity mismatch effect: performance degrades when the teacher’s capability diverges too far from the student’s.
+When using the large Qwen2-57B teacher, performance drops because the teacher is too strong relative to the student.
+With its vast world knowledge, the 57B model can often predict complex tokens purely from textual context, producing very low reference loss.
+However, the 7B student lacks this prior knowledge and instead depends on visual grounding to infer these tokens.
+As a result, the large teacher mistakenly signals ReGATE to prune these positions, causing the student to skip crucial computations.
+Conversely, the smaller Qwen2-1.5B teacher underestimates the student’s ability.
+It assigns high loss to linguistic patterns or factual details that the 7B student already handles easily.
+This floods the top- selection with linguistically difficult but visually irrelevant tokens, wasting computation on trivial cases instead of focusing on visual reasoning.
+Overall, these findings suggest that the optimal reference teacher should be capacity-aligned with the student.
+The reference loss is most effective when it reflects the student’s own uncertainty, directing computation precisely where textual priors fall short.
 | Benchmark | Response formatting prompts |
@@
 | ScienceQA | Answer with the option’s letter from the given choices directly. |
-| SEED-Bench | Answer with the option’s letter from the given choices directly. |
+| SEED | Answer with the option’s letter from the given choices directly. |
 | MLVU | – |
 | MVBench | Only give the best option. |
-| VideoMME | Answer with the option’s letter from the given choices directly. |
+| Video-MME | Answer with the option’s letter from the given choices directly. |
 | EgoSchema | Answer with the option’s letter from the given choices directly. |
@@
 | LongVideoBench | Answer with the option’s letter from the given choices directly. |
-Appendix C Additional Benchmarks Details
-Table 6 lists the evaluation prompts corresponding to each benchmark used in the experiments.
+8 Additional Benchmarks Details
+Table 9 lists the evaluation prompts corresponding to each benchmark used in the experiments, most of which are adapted from LMMs-Eval [zhang2024lmmseval].
+9 Qualitative Analysis of Reference Loss
+To validate the core mechanism of ReGATE, we qualitatively analyze the reference loss signal that guides its token selection.
+We assume that a high loss score from the text-only teacher indicates that a token requires visual information to be understood.
+Figure 5 shows two video Q&A examples, visualizing the loss for each word in the answer as calculated by a Mistral-7B [jiang2023mistral7b] teacher model.
+The results strongly support our assumption.
+As illustrated in the figure, tokens for visual details that are hard to guess from text alone, like the action “mixing” or the attribute “reflective”, get high loss scores.
+In contrast, simple grammatical words like “The” and “is”, or terms repeated from the question like “bartender”, get low scores.
+This difference confirms that reference loss is a reliable indicator of visual importance, enabling ReGATE to focus its computation on the most critical tokens for more efficient training.